OpenStudy (anonymous):

If sequence an (aka a sub n) is defined as an=sqrt(2+a(n-1)), and a1=sqrt(2), prove by induction that an is increasing.

ganeshie8 (ganeshie8):

@ikram002p

ganeshie8 (ganeshie8):

\(\large a_n^2 = 2 + a_{n-1}\)

OpenStudy (ikram002p):

wait why u need induction to show that its increasing ? just show that a_n <a_n+1 which is the same \((a_n)^2<a_{n+1}^2\)

OpenStudy (ikram002p):

ok lets induction since no response :P

OpenStudy (anonymous):

if you show that for the general case, that is essentially induction.

OpenStudy (ikram002p):

for first show for \((a_1)^2 =2\) \((a_2)^2=2+\sqrt2\) \(a_1^2<a_2^2 \) hold then assume its true for k \(a_k^2<a_{k+1}^2\) show for \(a_{k+1}^2<a_{k+2}^2\) ugh dont feel like continue :3 but its smothly could be done bbye

OpenStudy (anonymous):

It looks good, but I already got up to that part.

OpenStudy (anonymous):

the next step is what I was stuck on

ganeshie8 (ganeshie8):

\(\large\begin{align} a_{k+2}^2 &= 2 +a_{k+1} \\&\gt 2+a_k \\&=a_{k+1}^2 \end{align}\)

ganeshie8 (ganeshie8):

So, if \(\large a_k^2 \lt a_{k+1}^2\) holds, so does \(\large a_{k+1}^2 \lt a_{k+2}^2\) that ends the induction proof

OpenStudy (anonymous):

thanks